### v1

# Gravitational wave echos from physical black holes

Abstract
Gravitational wave echos from the coalescence of black hole binaries are often viewed as signals beyond general relativity or standard model. In this work, we show that these echos are inevitable in the black holes coalescence described by standard general relativity. This is because it is the physical black holes formed through gravitational collapse serve as the true description of astronomical black holes. For physical black holes, only their asymptotic structure before the horizon forms are detectible to the outside probes. Here, by investigating the scattering of a gravitational wave burst on a physical black hole and pay special attention to the echos in the waveform, we uncover distinct features of the echos both in the time and frequency domains.
$^{1}$School of Physics and Optoelectronic Engineering, Beijing University of Technology, Beijing, China
$^{2}$School of Physics and Astronomy, Yunnan University, Kunming 650091, PR China
1 Introduction
It is commonly accepted that black holes with central singularity or circular-line-like singularity are the fate of massive bodies’ evolution under their self gravitation. So static or stable solutions to the Einstein equation such as the Schwarszchild and Kerr-type are taken as representation of black holes in astronomy for granted. However, serious reflection suggests that the situation is more subtle than commonly believed [1, 2, 3, 4, 5, 6, 7, 8, 9]. This subtlety arises from the time definition dependence of physics in General relativity. More precisely for black holes, it is a matter of “when” does the gravitational collapse complete.
One can see for example in the Oppenheimer-Snyder model [10, 11], all probes living outside the collapsing material will see the surface of the collapsar undergoes an endless contraction towards the event horizon but never reach it at all. To these probes, detectable physical events happen in the Schwarzschild time definition, during which the collapse never finishes, so the horizon never forms in any finite future. While to probes co-moving with the collapsing material, although they will see the surface of the collapsar fall across the event horizon and hit on the singularity in finite duration, what they detect after the horizon formation happens in the domain of Lemaitre time definition but beyond that of the Schwarzschild time definition, which is inaccessible to outside probes. To all outside probes, the physics concerning them happen in the Schwarzschild time definition, and the spacetime has an endless imploding and horizonless structure. To emphasize this fact we refer black holes formed through gravitational collapse as “physical black holes”, while those described by the Schwarzschild or Kerr-metric as “mathematical black holes”. Although many works concerning the black hole phenomenology use mathematic black holes where in fact physical black hole is expected, some authors distinguish this two concepts carefully and emphasize the necessity in both observational [1, 2] and theoretical aspects [3, 4, 5].
Since the first gravitational wave signal detected by LIGO ten years ago [12], the world witnessed a surge of growth in gravitational wave astronomy. Now, about a hundred of gravitational wave events have been reported [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], most of which are from black hole binary coalescence. With the rapid growth of the gravitational wave antenna sensitivity and the data processing efficiency, in the future, more detailed structures of the waveform may be detected, among which the gravitational wave echo from near horizon structure of the black holes is a possibility.
Gravitational wave echos from the binary coalescence waveform are widely viewed as a smoking gun for new physics such as quantum black holes [24, 25, 26, 27, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], exotic compact objects (ECOs) [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], black hole remnants [50], modified gravity theory [51, 52, 53, 54, 55, 68], dark matter cloud [57, 58], wormhole [59, 60, 61, 62] and etc. Experimentally, possible echo signals have been reported in black hole coalescence events GW150914, GW151226, and LVT151012 [63, 64], while their authenticity are under debate [65, 66, 67, 68, 69]. Currently, more powerful methods are under development aiming at digging echos from the recorded black hole coalescence events [70]. Theoretically, the echos can be attributed to the boundary condition for gravitational perturbations near the event horizon. In some phenomenological works on ECOs, the Dirichlet boundary condition is applied at the ECO surface for simplicity [24]. But to get precise signals, the penetration of gravitational wave into the matter region need be considered, on this issue the readers are referred to [49, 71, 72, 73, 74, 75, 76].
In this work, we focus on the gravitational wave echos from a gravitational burst scattering off a neutral and non-spinning physical black hole. Of course, our method used here has no conceptual difficulty generalizing to the spinning cases. Neglecting the angular momentum is a purely technique simplicity motivated choice. For this black hole, the endless contracting surface serves as a moving reflector modeled by a Dirichlet boundary condition. We observe that the echos from this black hole are fainter than those from ECOs. Their most distinctive feature lies in their frequency content, that is, the huge and dynamical Doppler red shift between the neighboring echos. And the time intervals between neighboring echos from are un-equal, they grow as the time increases.
Our paper is organized as follows. In Sec. II we review the definition of physical black holes and the dynamics of its gravitational perturbation. In Sec. III we go through the details of wave packet scattering of physical black holes, where the waveform is presented along with analytical and numerical analysis. Conclusions and discussions are provided in Sec. IV.
Throughout this work the units will be chosen in such a way that $G_{N}=c=1$.
2 Physical black holes
Consider a neutral and non-spinning collapsar with mass $M$. In the $\{t,r\}$ plane of the spacetime diagram, the surface of the star contracts following time-like world lines which asymptotically approach the event horizon so that
$\frac{dt}{dr}<-\frac{1}{1-\frac{2M}{r}}.$ (1)
At the very late stage of the gravitation collapse the star surface locates on $r_{\varepsilon}(t)=2M+\varepsilon(t)$, where $\varepsilon(t)$ is a small parameter.
Since the self-gravitation dominates all other interactions and the degeneration pressure during this stage of the collapse, researchers choose neglecting the pressure of the collapsing material and approximating them as pressureless dust, [1, 2, 11, 77, 78]. In this case, the motion of the collapsar surface is freely falling. In this work, we will let the collapsar surface follow an arbitrary ingoing time-like trajectory for generality. The spacetime metric outside the surface is still the Schwartzchild metric while that inside is the solution of Einstein equation supported by energy momentum tensors contributed by the collapsing materials.
At the ring-down phase of a binary merger process, the spacetime geometry can be decomposed into a background physical black hole plus a small perturbation. The later can be expanded into superpositions of many spherical harmonics whose radial part is governed by the master equation [24]
$\frac{d^{2}}{dt^{2}}-\frac{d^{2}}{dx^{2}}+V_{l}(r)]R_{l}(t,x)=0,$ (2)
where $x=r+2M\ln(\frac{r}{2M}-1)$ is the tortoise coordinate and
$V_{l}(r)=(1-\frac{2M}{r})[\frac{l(l+1)}{r^{2}}+\frac{(1-s^{2})2M}{r^{3}}]$ (3)
is the centrifuge potential peaked at the photon sphere radius $r=3M$. Here we take $s=2$ for the gravitational wave mode perturbation. For scalar and vector mode perturbations one can simply set $s=0,1$. In the following we will concentrate on the quadrupolar perturbation $l=2$ and write $R_{2}(t,x)$ as $\psi(t,x)$ in the sequel.
To solve the field equation (2) we need to specify the boundary conditions at infinity and on the collapsar surface $r_{\varepsilon}(t)$. The former will be taken as the purely radiational or Sommerfeld condition $(\partial_{t}+\partial_{x})\psi=0$, while the later is assumed to be Dirichlet boundary condition
$\psi(t,X_{\varepsilon}(t))=0,$ (4)
where $X_{\varepsilon}(t)$ is the trajectory of the collapsar surface expressed in the tortoise coordinate.
3 Echos from the scattering of wave packets
3.1 Analytical estimations
In the following we will consider the scattering of a Gaussian shaped gravitational wave burst. For estimation we neglect the width of the centrifuge potential and the burst itself at first. So the propagation of the wave burst can be studied with geometry optics. Note that the gravitational wave travels at the speed of light so it will surely catch up with the collapsar surface and get bounced back, as sketched in Fig.(1).
Figure 1: The spacetime diagram of a gravitational wave’s scattering on a physical black hole. The left solid curve represents the star surface while the right vertical dashed line represents the photon sphere.
3.1.1 time-delay
Now we consider an ingoing pulse hitting on the photon sphere at $t=0$. A portion of the pulse will be reflected by the centrifuge potential and travel outward along a light-like geodesic. To all probes or observers living far away $R\gg 3M$, the estimated time at which the first signal will be recorded at is
$\begin{split}t_{1}&=\int_{3M}^{R}dr\frac{1}{1-\frac{2M}{r}}\\ &=(R-3M)+2M\ln\frac{R-2M}{M}.\end{split}$ (5)
The transmitted part of the pulse will keep going inward until it hit the collapsar surface and get bounced back. Part of this signal will penetrate the centrifuge potential and recorded by the observer as the first echo at time
$\begin{split}t_{2}&=t_{1}+2\int_{3M}^{2M+\varepsilon_{1}}dr\frac{1}{1-\frac{2M}{r}}\\ &=t_{1}+2(M+2M\ln\frac{M}{\varepsilon_{1}}),\end{split}$ (6)
where $\varepsilon_{1}$ is the distance between the Schwartchild radius and the collapsar surface when it is hit on by this pulse, as shown in Figure 1.
For two neighboring echos, the difference between their recorded time is
$2(\int_{3M}^{2M+\varepsilon_{n-1}}dr\frac{1}{1-\frac{2M}{r}}-\int_{3M}^{2M+\varepsilon_{n}}dr\frac{1}{1-\frac{2M}{r}})=2(\varepsilon_{n}-\varepsilon_{n-1}+2M\ln\frac{\varepsilon_{n}}{\varepsilon_{n-1}})$ (7)
Since the physical black hole has already entered its very late time of the collapse, $\varepsilon_{i}\to 0,i=1,2,\cdots$. In this case Eq.(7) can be approximated by $4M\ln(\frac{\varepsilon_{n}}{\varepsilon_{n-1}})\sim-4M\frac{\varepsilon_{n-1}-\varepsilon_{n}}{\varepsilon_{n-1}}$.
3.1.2 Redshift
We now assume that the wave burst is monochrome. In the Schwartzchild coordinate, its wave vector can be written as
$K^{\mu}=\{\frac{\omega}{1-\frac{2M}{r}},\omega,0,0\}$ (8)
where $\omega$ is the observed frequency by static observers located on $r$. If the world line of the incoming wave and collapsar surface intersect at $(t,r)$, then to the co-moving observer fixed on the collapsar surface, the frequency would be
$\omega^{\prime}=-g_{ab}K^{a}(\frac{\partial}{\partial\tau})^{b}=(\dot{t}+\frac{\dot{r}}{1-\frac{2M}{r}})\omega.$ (9)
The over dot here denotes differentiations with respect to the collapsing surface’s proper time. After reflection the outgoing wave vector in the Schwartzchild coordinate can be written in the form
$K^{\prime u}=\{\frac{\omega^{\prime\prime}}{1-\frac{2M}{r}},-\omega^{\prime\prime},0,0\}$ (10)
By the relation $\omega^{\prime}=-g_{ab}K^{\prime a}(\frac{\partial}{\partial\tau})^{b}$, we know
$\omega^{\prime\prime}=\frac{\omega^{\prime}}{\dot{t}+\frac{\dot{r}}{1-\frac{2M}{r}}}=\frac{\dot{t}+\frac{\dot{r}}{1-\frac{2M}{r}}}{\dot{t}-\frac{\dot{r}}{1-\frac{2M}{r}}}\omega.$ (11)
From Eq.(11) we see that there is a huge redshift in the reflected signal if the speed of the collapsar surface reaches $\mathrm{O}(1)$ fraction of the local speed of light.
3.2 The waveform and echos
3.2.1 Setups
To obtain the gravitational waveform, we will numerically simulate the propagation of wave bursts according Eq.(2) with the aforementioned boundary condition. In our simulations, the mass is normalized to $M=1$, and the spatial infinity is placed at $x=100$. In mathematics, we known that the moving Dirichelet boundary condition (4) is indistinguishable from a delta function barrier with infinite height placed in the same place. In our program, we use a thin Gaussian function shaped potential to do this work. So the total potential in the master equation Eq.(2) becomes
$V(x)=V_{l=2}(r)+V_{B}(x,X(t)).$ (12)
The boundary potential $V_{B}(x,X(t))=he^{-\frac{(x-X(t))^{2}}{d}}$ here is the thin Gaussian barrier used to model the collapsar surface and the parameters are chosen as $h=100$ and $d=1$, respectively.
The initial waveform is chosen as
$\psi(0,x)=e^{-\frac{(x-x_{0})^{2}}{\sigma^{2}}},\partial_{t}\psi(0,x)=0,$ (13)
which is centered at $x_{0}=10$ and broadened to width $\sigma=2$. What we are going to plot is the waveform recorded by observers placed at $x=70$. The motion of the collapsar surface is chosen as
$X_{\varepsilon}(t)=-30-\frac{t}{4},$ (14)
corresponding to the collapsar surface contracting at $1/4$ times of the speed of light. The speed of the collapsar surface depends on the collapsing material’s equation of state, so is not uniquely determined. Our choice here is a result of taking balance between the need of clear demonstration and saving of computational resources.
3.2.2 Waveform
The projected waveform is presented in Figure 2, where the duration before the echos is discarded. The portion of the initial wave burst that penetrated the centrifuge potential will catch up with the collapsar surface and be reflected. After that it will be scattered by the centrifuge potential again. Part of this signal will transmit through the potential and gives rise to the first echo in the waveform recorded by the gravitational wave detector at $t\sim 150-240$. The other part of the signal scattered back by the centrifuge potential will repeat the same process and form a series of echos. In Figure 2, we can see the second echo around $t\sim 290-380$.
Figure 2: The echos from a Gaussian wavepacket scattering off a physical black hole. Two echos can be seen in this waveform, one around 150240150-240 and another one around 290380290-380.
Figure 3: The echos from a Gaussian wavepacket scattering off an ECO. Three echos can be seen in this waveform, the first lies around 190300190-300, the second lies around 315425315-425 and the last one starting from $440$. The three echos are equally spaced with interval of $125$.
We put the waveform of echos from an ECO [24, 42, 43] in Figure 3 for comparison, where all other parameters are chosen identically as above except for the star surface location. For the ECO in Figure 3, the star surface is statically located at x=60x=-60. The figure shows three distinct echo signals, the first two are roughly located at $t\sim 190-300$ and $t\sim 315-425$, while the third starts around $t\sim 440$ and persists till the end of the simulation.
Comparing the two waveforms we immediately see two differences. The first is that echos from physical black holes are not equally spaced, while those from ECO are equally spaced. Among the three echos in Figure 3 we see that the two intervals between neighboring 3 echos are both roughly $125$ . In Figure 2, the interval between the first two echos is about 140. The third echo should appear at around $t\sim 440$ if the echos are equally spaced. However, it does not manifest till the end of the simulation at $t=450$. This confirms our estimation in Sec.3.1.1. The second difference is the strength of the echos. We can see that the first echo in Figure 2 seems slightly weaker than its sibling in Figure 3 while the second echo in Figure 2 is clearly weaker than its counterpart in Figure 3. The reason behind this phenomenon is as follows. The ingoing wave receives Doppler red shift when reflected by the collapsar surface. When it is scattered by the centrifuge potential, which tends to reflect the low frequency components back. Thus the echos from the physical black holes are fainter than those from ECO. In the next part we will dive deeper on this issue.
3.2.3 Spectra
This part applies Fourier analysis to echos in Figures 2 and 3 to examine their frequency spectrum. We see that each of the echos are isolated inside a specific time interval. Hence the Fourier transformation of the echos can be computed via [24]
$\psi(\omega)=\int_{t_{i}}^{t_{f}}dt\psi(t)e^{i\omega t},$ (15)
where $t_{i}$ and $t_{f}$ are the beginning and the ending of the time interval of a specific echo signal, respectively. The spectrum of echos from physical black holes are presented in Fig.4(a) and (b), while that from ECO are presented in Fig.4(c) and (d). We see that the peak frequency of the first and second echos from ECO are both located at roughly $0.4$. For the physical black hole, the peak frequency of the first echo is about $0.4$, while that of the second echo is about $0.25$. This is a quantitative demonstration of Doppler shift from the reflection on the collapsar surface. Comparing Figure 4(b) and 4(d), we easily see that the second echo from physical black holes is fainter than that from ECO. Examine Figure 4(a) and 4(c), since the area enclosed by the power spectrum and horizontal axis represents the energy of the echo signal, we can confirm that the first echo from physical black hole is also weaker than that from ECO, as expected. The reason this difference is not so clear in the time domain waveform Figure 2 and 3 is that, the radial speed of the collapsar surface is set to be relatively small in order to see desired echo signal in our computational resources, thus yielding a smaller Doppler red shift factor.
Figure 4: Subdiagrams (a) and (b) display the absolute value of the Fourier spectrum of the first echo and second echo from physical black holes. Those from ECO are displayed in subdiagrams (c) and (d) respectively.
Figure 4(c) and (d) give a demonstration of our statements on the relation between the reflectivity of the centrifuge potential and wave frequencies at the end of above subsection. Comparing Figure 4(c) and (d) we see that the high frequency component in the first echo is absent in the second echo. Similar things also happen in Figure 4 (a) and (b) but not so obviously. This can be understood the following way. We assign a reflection factor $R(\omega)$ and a transmitting factor $T(\omega)=1-R(\omega)$ to the centrifuge potential. For a specific potential with finite height, $T(\omega)$ has an overall dentation to decrease to the lower frequency region. This explains what we see in Figure 4(c) and (d). That is, the high frequency component of the first echo is transmitted outside through the centrifuge potential and the low frequency component is reflected back and left for the second echo.
4 Conclusion and discussion
Astrophysical black holes are formed through gravitational collapse. The time dilation mechanism of general relativity implies that in the Schwartzchild time definition, the entire life of a collapsar is lived in dynamical and horizonless geometry eternally approaches to the formation of the event horizon but never to reach it. The late stage of the collapsars are referred as physical black holes. The reason behind the privilege of Schwartzchild time in black hole phenomenology lies in that all observable signal comes from probes living outside the black holes, whose entire history lie only in the domain of Schwartzchild time definition. With the aim of establishing a testable signal to distinguish the physical black holes from mathematical ones, we considered the gravitational wave echos from a gravitational burst scattering off a physical black hole. We found that such echos are inevitable consequences for physical black holes and have distinct features comparing with those discussed in other scenarios. The differences can be seen in both time and frequency domain. Our results indicate that the presence of gravitational wave echos does not necessarily implies the existence of new physics. Rather, it is comprehensible within the framework of classical general relativity.
One may wonder since the physical black hole and mathematical black holes are so different, why did’t researchers rule one of them out by existing observations? The answer is somewhat surprising. Although astronomers have registered hundreds of black holes from the binary merger process, the existence of event horizon is still yet to be confirmed [79]. On this issue we recommend the readers to refs.[79, 80], where it is pointed out that the ring-down signals are mainly determined by physics of the photon sphere, instead of the physics of event horizon. With identical hair parameters, physical black holes, mathematical black holes and ECO black hole mimickers all have similar ring-down waveforms because all of them have the same photon sphere structure. Our result suggests that the gravitational wave echo is a potential signal to distinguish them from each other.
The original signal stimulating echos arises from the ring-down motion of a binary merger system. A natural question arises here, are there echos produced druing the inspiral and merger motion of the binary system? The answer is yes. During all phases, the gravitational wave propagates all directions. Most of it will radiate freely to infinity while a small fraction will hit the binary system surfaces and bounce back, thus forming echos. It is noteworthy that the feature of echos from these phases could be drastically different from our results even for binary systems composed of two non-spinning black holes. The key reason is, during the inspiral and merger phases, the matter surface providing the reflective condition is far from spherically symmetric.
Acknowledgements
This research was funded by the National Natural Science Foundation of China under Grant No.11785082.